# On the eigenvalue problem for a bulk/surface elliptic system

### v1

@@ 2. Preliminaries
In the case A) above both $\Gamma_{0}$ and $\Gamma_{1}$ are relatively open and compact, so we can repeat previous arguments and get (28) and (29) with $\Gamma$ replaced by $\Gamma_{0}$ and $\Gamma_{1}$. Hence (28) and (29) continue to hold when replacing $\Gamma$ with $\Gamma_{0}$ and $\Gamma_{1}$, so we can set $\Delta_{\Gamma_{i}}\in\mathcal{L}(W^{1,p}(\Gamma_{i});W^{-1,p}(\Gamma_{i}))$, $i=0,1$, by replacing $\Gamma$ with $\Gamma_{i}$ in (30). In this way we get
@@
Moreover, using the splitting (26) one obtains $\Delta_{\Gamma}=(\Delta_{\Gamma_{0}},\Delta_{\Gamma_{1}})$, so by the identification (27) one gets $\Delta_{\Gamma_{1}}={\Delta_{\Gamma}}_{|W^{s+1,p}(\Gamma_{1})}$. Hence the symbols $\Delta_{\Gamma_{1}}$ and $\Delta_{\Gamma}$ can be used with the same meaning in the present paper.
@@
The result is well–known when $\Gamma$ is smooth, see for example [43, p. 28]. Moreover the proofs in the cases $p=2$ and $s=1$ have been already explicitly given, respectively see [36, Theorem 5.0.1] and [47, Lemma B.1, Appendix B]. Since in the present paper we shall use only these cases, here we just sketch how to generalize the arguments used in the proof of [47, Lemma B.1, Appendix B] from $|s|\leq 1$ to $|s|\leq r-1$ (this fact being nontrivial only when $r>2$).
When $s\in[1,r-1]$, using the standard localization technique exactly as in the proof of [47, Lemma B.2, Appendix B], as well as higher elliptic regularity, see for example [21, Chapter 2, Theorem 2.5.1.1., p. 128], one gets that $-\Delta_{\Gamma_{1}}u+u\in W^{s-1,p}(\Gamma_{1})$ implies $u\in W^{s+1,p}(\Gamma_{1})$, so $A_{s,p}$ is surjective. Being injective when $s=1$ one then obtains that $A_{s,p}$ is bijective for $s\in[0,r-1]$. The proof is then completed by transposition. ∎
@@ 2.4. The trace and normal derivative operators
By [21, Chapter 1, Theorem 1.5.1.2, p.38], the standard trace operator $u\mapsto u_{|\Gamma}$ from $C(\overline{\Omega})$ to $C(\Gamma)$, when restricted to $C(\overline{\Omega})\cap W^{m,p}(\Omega)$, $1<p<\infty$, $m\in\mathbb{N}$, $1\leq m\leq r$, has a unique surjective extension $\operatorname{Tr}_{\Gamma}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma)\right)$. Moreover, when $m=1$, $\operatorname{Tr}_{\Gamma}$ has a bounded right–inverse, i.e., $\mathcal{R}\in{\mathcal{L}}\left(W^{1-\frac{1}{p},p}(\Gamma),W^{1,p}(\Omega)\right)$ such that $\operatorname{Tr}_{\Gamma}\cdot\mathcal{R}=I$, independent on $p$. We shall denote, as in § 1, $\operatorname{Tr}_{\Gamma}u=u_{|\Gamma}$, for simplicity.
Moreover, denoting by $u_{|\Gamma_{i}}$, for $i=0,1$, the restriction of $u_{|\Gamma}$ to $\Gamma_{i}$, and $\operatorname{Tr}_{\Gamma_{i}}u=u_{|\Gamma_{i}}$, one clearly has $\operatorname{Tr}_{\Gamma_{1}}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma_{1})\right)$. When $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}=\emptyset$ one also gets $\operatorname{Tr}_{\Gamma_{0}}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma_{0})\right)$.
Moreover, when $r\geq 2$, i.e. $\Gamma$ is $C^{2}$, for all $p\in(1,\infty)$, $m\in\mathbb{N}$ such that $2\leq m\leq r$, and $u\in W^{m,p}(\Omega)$, denoting $\partial_{\nu}u=\sum_{i=1}^{N}\partial_{i}u_{|\Gamma}\,\nu^{i}$, where $\nu=(\nu^{1},\ldots,\nu^{N})$, and $\partial_{i}u_{|\Gamma_{i}}$ are taken in the trace sense sense, we get the normal derivative operator
@@
Clearly $\partial_{\nu}u$, which restrictions to $\Gamma_{i}$ will be denoted as $\partial_{\nu}u_{|\Gamma_{i}}$, can be defined in such a trace sense only when $r\geq 2$ and $u\in W^{2,p}(\Omega)$. Hence, when $r=1$, we set the normal derivative in a distributional sense as follows.
@@
We recall, see [45, Lemma 1, p. 2147] which trivially extends to $\Gamma$ of class $C^{1}$, that the space
@@
On the other hand, by the assumption $\mathcal{H}^{N-1}(\Gamma_{0})>0$, the norm $\|\cdot\|_{H^{1}}$ given by (9) is equivalent to ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{H^{1}}$, as it is well–known, see for example [50, Lemma 1, p. 8]. Hence in the paper, as stated in § 1, we shall endow the space $H^{1}$ with $\|\cdot\|_{H^{1}}$.
@@ 3. Analysis of problem (13)
This section is devoted to the preliminary analysis of problem (13). We start by making precise which types of solutions we are going to consider.
@@
    When assumption (R) holds, $u\in H^{2}$ is a strong solution of (13) provided
@@
Essentially as in [47], it is useful to deal with weak solutions of (13) at a more abstract level. By [47, Lemma 2.1], extended to the complex case, the trivial embedding $H^{1}\hookrightarrow H^{0}$ is dense. In the sequel we shal identify $L^{2}(\Omega)$ and $L^{2}(\Gamma_{1})$ with their duals $[L^{2}(\Omega)]^{\prime}$ and $[L^{2}(\Gamma_{1})]^{\prime}$ coherently to the distributional identification. We shall also identify $L^{2}(\Omega)$ and $L^{2}(\Gamma_{1})$ with their isometric copies $L^{2}(\Omega)\times\{0\}$ and $\{0\}\times L^{2}(\Gamma_{1})$ contained in $H^{0}=L^{2}(\Omega)\times L^{2}(\Gamma_{1})$. Consequently we shall identify $H^{0}$ with its dual $(H^{0})^{\prime}$ according to the identity
@@
We can then introduce the chain of dense embeddings, or Gel’fand triple,
@@
(41) $\langle A_{1}u,v\rangle_{H^{1}}=(u,v)_{H^{1}}=\int_{\Omega}\nabla u\nabla\overline{v}+\int_{\Gamma_{1}}(\nabla_{\Gamma}u,\nabla_{\Gamma}v)_{\Gamma},$
as well as its part $A:D(A)\subset H^{1}\subset H^{0}\to H^{0}$ given by
@@
were (39) is used. By a quick comparison between (37) and (41)–(42) one gets the following result.
@@
Combining it with (43) we get (37), that is $u$ is a weak solution of (13). By Lemma 8 thus $u\in D(A)$, so $H^{2}\subseteq D(A)$.
@@
Before stating our next result we remark that, since the operator $A$ is trivially closed, we can endow $D(A)$ with the graph norm $\|\cdot\|_{D(A)}=\|\cdot\|_{H^{1}}+\|A(\cdot)\|_{H^{0}}$, obtaining a Hilbert space. The following result shows that problem (13) is well–posed and that $A$ is an isomorphism.
@@
    If $f\in L^{p}(\Omega)$ and $g\in L^{p}(\Gamma_{1})$ for $p\in[2,\infty)$ then $u\in W^{2,p}$. Moreover, there is a positive constant $c_{2}=c_{2}(\Omega,\Gamma_{1},p)$ such that
@@
    (46) $\|u\|_{H^{m}}\leq c_{3}\left(\|f\|_{H^{m-2}(\Omega}+\|g\|_{H^{m-2}(\Gamma_{1})}\right)$
@@
We start by proving the first assertion in ii), by induction on $m$. When $m=2$, by Lemma 9 we have $u\in H^{2}$. Now we suppose by the induction hypotesis that $m\geq 3$, the assertion holds for $m-1$, $f\in H^{m-2}(\Omega)$ and $g\in H^{m-2}(\Gamma_{1})$. We claim that $u\in H^{m}$. By the induction hypotesis we have $u\in H^{m-1}$. Consequently, by (33), $\partial_{\nu}u_{|\Gamma_{1}}\in H^{m-5/2}(\Gamma_{1})$. Since, by Lemma 9, we have $\partial_{\nu}u_{|\Gamma_{1}}-\Delta_{\Gamma_{1}}u=g$ on $\Gamma_{1}$, by Lemma 6 we get $u_{|\Gamma_{1}}\in H^{m-1/2}(\Gamma_{1})$. By the same Lemma we also have $-\Delta u=f$ in $\Omega$, so by higher elliptic regularity (see [21, Chapter 2, Theorem 2.5.1.1, p. 128]) we get $u\in H^{m}(\Omega)$. A further application of (33) then yields $\partial_{\nu}u_{|\Gamma_{1}}\in H^{m-3/2}(\Gamma_{1})$. Since $\partial_{\nu}u_{|\Gamma_{1}}-\Delta_{\Gamma_{1}}u=g$ on $\Gamma_{1}$, again by Lemma 6 we get $u_{|\Gamma_{1}}\in H^{m}(\Gamma_{1})$, proving our claim.
@@
It is then clear that, in this way, the following alternative occurs: either after finitely many iterations we get $\partial_{\nu}u_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$, so the proof of our claim can be completed as when $N=2$, or we can go on indefinitely. In this case the recursive formula $p_{0}=2$, $p_{n+1}=\frac{(N-1)p_{n}}{N-p_{n}}$, defines a strictly increasing sequence $(p_{n})_{n}$ such that $1<p_{n}<\min\{p,N\}$ for all $n\in\mathbb{N}$. Let $l:=\lim_{n}p_{n}\in(1,\min\{p,N\}]$. Passing to the limit in the recursive formula we get that $l=N$ leads to $l=\infty$, a contradiction, so $l<N$. Passing to the limit then $l=\frac{(N-1)l}{N-l}$, that is $l=1$, a contradiction. Hence after finitely many iterations we get $\partial_{\nu}u_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ and conclude the proof of our claim. To prove (45) and conclude the proof one then applies the same argument already used to prove (46). ∎
@@ 4. Proofs
This section is devoted to prove our main results, already stated in Section 1.
@@
We are then going to apply to $B$ the standard Spectral Decomposition Theorem for compact and self adjoint operators, see for example [6] or [14]. The spectrum $\sigma(B)$ of $B$ then consists of 00 and of its point spectrum $\sigma_{p}(B)\subset\mathbb{R}$, which is at most countable. Moreover, since $A$ is bijective, $0\not\in\sigma_{p}(B)$ and, by the Fredholm Alternative, $\text{ker}(B-\mu I)$ is finite–dimensional for all $\mu\in\sigma_{p}(B)$. Moreover $H^{0}$, being separable and infinite dimensional, admits a Hilbert basis of eigenvectors of $B$. Hence $\sigma_{p}(B)=\{\mu_{n},n\in\mathbb{N}\}$, where the sequence $(\mu_{n})_{n}$ is injective and $\mu_{n}\to 0$ as $n\to\infty$. Consequently $\Lambda=\{\lambda_{n},n\in\mathbb{N}\}$, where $\lambda_{n}=1/\mu_{n}$. Let us denote $H_{n}=\text{ker}(B-\mu_{n}I)$. Since, for each $u\in H_{n}$, we have $\lambda_{n}=\|u\|_{H^{1}}^{2}$, it follows that $\Lambda\subset(0,\infty)$. Consequently we can re–arrange the $\lambda_{n}$’s, according to their finite multiplicities, as in (11).
@@
At first we prove i), by combining Theorem 11–i) with a standard bootstrap argument. More in detail, by Lemma 9, $u_{n}\in H^{2}$ for each $n\in\mathbb{N}$, that is $u_{n}\in H^{2}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in H^{2}(\Gamma_{1})$. We now distinguish between the cases $N\leq 4$ and $N\geq 5$. In the first one, by applying Morrey’s Theorem when $N<4$ and the limiting case of Sobolev Embedding Theorem when $N=4$, we get $u_{n}\in L^{p}(\Omega)$ for all $p\in[2,\infty)$. By applying the same results on $\Gamma_{1}$ we get that ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ for all $p\in[2,\infty)$ provided $N\leq 5$, so a – fortiori when $N\leq 4$. By applying Theorem 11–i) we then get $u_{n}\in W^{2,p}$ for all $p\in[2,\infty)$. When $N\geq 5$ we further distinguish between the case $5\leq N\leq 8$ and $N\geq 9$. In the first one, by Sobolev Embedding Theorem, we have $u_{n}\in L^{q_{1}}(\Omega)$, where $\frac{1}{q_{1}}=\frac{1}{2}-\frac{2}{N}$, and as above one also gets that $u_{n}\in L^{q_{1}}(\Gamma_{1})$. Hence, by Theorem 11–i), one has $u_{n}\in W^{2,q_{1}}$. Since $\frac{1}{q_{1}}-\frac{2}{N}=\frac{1}{2}-\frac{4}{N}\leq 0$, by applying Morrey’s Theorem when $N<8$ and the limiting case of Sobolev Embedding Theorem when $N=8$, we get $u_{n}\in L^{p}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ for all $p\in[2,\infty)$, so $u_{n}\in W^{2,p}$ for all $p\in[2,\infty)$ as in the previous case. It is then clear how to handle the remaining case $N\geq 9$, proving after finitely many iterations that $u_{n}\in L^{p}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ , so $u_{n}\in W^{2,p}$, for all $p\in[2,\infty)$. The conclusion $u_{n}\in C^{1}(\overline{\Omega})$ then follows from Morrey’s Theorem.
To prove ii) we remark that, since $u_{n}\in W^{2,p}$, for all $p\in[2,\infty)$, using (33) we have $\partial_{\nu}{u_{n}}_{|\Gamma_{1}}\in W^{1-\frac{1}{p},p}(\Gamma_{1})$ for all $p\in[2,\infty)$. We also remark that $W^{1-\frac{1}{p},p}(\Gamma_{1})$ coincides with the Besov space $B^{1-\frac{1}{p},p}_{p}(\Gamma_{1})$, as proved in [44, pp. 189–190]. By Morrey’s Theorem for Besov spaces, see [28, Chapter 17, Theorem 17.52, p. 562], taking $p>N$, we have $\partial_{\nu}{u_{n}}{|\Gamma_{1}}\in C^{0,1-\frac{N}{p}}(\Gamma_{1})$. Moreover, since we also have ${u_{n}}_{|\Gamma_{1}}\in C^{1}(\Gamma_{1})$, by (12) and Lemma 9, $-\Delta_{\Gamma_{1}}u_{n}=f_{n}\in C^{0,1-\frac{N}{p}}(\Gamma_{1})$. Being $\Gamma_{1}$ compact, we can use the standard localization technique together with the classical Schauder Estimates, see [6, Theorem 9.33] or [19, Chapter 6], to recognize that in local coordinates $u_{n}$ is of class $C^{2,1-\frac{N}{p}}$, so $u_{n}\in C^{2}(\Gamma_{1})$, proving ii).
@@
Moreover, since $\Gamma$ is $C^{2}$, it is also $C^{1,1}$, so $\Omega$ satisfies the interior ball condition, that is to say there is an open ball $B\subseteq\Omega$ such that $x_{0}\in\partial B$. See [29, Corollary 2,p. 550] or [3, Theorem 1.0.9,p. 7]. Now, since $u\in C^{\infty}(B)\cap C^{1}(\overline{B})$, $-\Delta u=\lambda_{1}u\geq 0$ in $B$, and $u(x)>u(x_{0})=0$ for all $x\in B$, we can apply the Höpf Lemma, see [17, Chapter 6,p. 330], to conclude that $\partial_{\nu}u(x_{0})<0$, contradicting (60) and completing the proof. ∎

### v2

@@ 2. Preliminaries
In the case A) defined above both $\Gamma_{0}$ and $\Gamma_{1}$ are relatively open and compact, so we can repeat previous arguments and get (28) and (29) with $\Gamma$ replaced by $\Gamma_{0}$ and $\Gamma_{1}$. Hence (28) and (29) continue to hold when replacing $\Gamma$ with $\Gamma_{0}$ and $\Gamma_{1}$, so we can set $\Delta_{\Gamma_{i}}\in\mathcal{L}(W^{1,p}(\Gamma_{i});W^{-1,p}(\Gamma_{i}))$, $i=0,1$, by replacing $\Gamma$ with $\Gamma_{i}$ in (30). In this way we get
@@
Moreover, using the splitting (26), one obtains $\Delta_{\Gamma}=(\Delta_{\Gamma_{0}},\Delta_{\Gamma_{1}})$, so by the identification (27) one gets $\Delta_{\Gamma_{1}}={\Delta_{\Gamma}}_{|W^{s+1,p}(\Gamma_{1})}$. Hence the symbols $\Delta_{\Gamma_{1}}$ and $\Delta_{\Gamma}$ can be used with the same meaning in the present paper.
@@
The result is well–known when $\Gamma$ is smooth, see for example [43, p. 28]. Moreover the proofs in the cases $p=2$ and $s=1$ have been already explicitly given, respectively see [36, Theorem 5.0.1] and [47, Lemma B.1, Appendix B]. Since in the present paper we shall use only these cases, here we just sketch how to generalize the arguments used in the proof of [47, Lemma B.1, Appendix B] from $|s|\leq 1$ to $|s|\leq r-1$ (this generalization being needed only when $r>2$).
When $s\in[1,r-1]$, using the standard localization technique exactly as in the proof of [47, Lemma B.2, Appendix B], as well as elliptic higher regularity results, see for example [21, Chapter 2, Theorem 2.5.1.1., p. 128], one gets that $-\Delta_{\Gamma_{1}}u+u\in W^{s-1,p}(\Gamma_{1})$ implies $u\in W^{s+1,p}(\Gamma_{1})$, so $A_{s,p}$ is surjective. Being injective when $s=1$ one then obtains that $A_{s,p}$ is bijective for $s\in[0,r-1]$. The proof can then be completed by transposition. ∎
@@ 2.4. The trace and normal derivative operators
By [21, Chapter 1, Theorem 1.5.1.2, p. 38], the standard trace operator $u\mapsto u_{|\Gamma}$ from $C(\overline{\Omega})$ to $C(\Gamma)$, when restricted to $C(\overline{\Omega})\cap W^{m,p}(\Omega)$, $1<p<\infty$, $m\in\mathbb{N}$, $1\leq m\leq r$, has a unique surjective extension $\operatorname{Tr}_{\Gamma}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma)\right)$. Moreover, when $m=1$, the operator $\operatorname{Tr}_{\Gamma}$ has a bounded right–inverse, i.e. an operator $\mathcal{R}\in{\mathcal{L}}\left(W^{1-\frac{1}{p},p}(\Gamma),W^{1,p}(\Omega)\right)$ such that $\operatorname{Tr}_{\Gamma}\cdot\mathcal{R}=I$, with $\mathcal{R}u$ independent on $p$. For the sake of simplicity we shall denote, as in § 1, $\operatorname{Tr}_{\Gamma}u=u_{|\Gamma}$.
Moreover, denoting by $u_{|\Gamma_{i}}$, for $i=0,1$, the restriction of $u_{|\Gamma}$ to $\Gamma_{i}$, and also denoting $\operatorname{Tr}_{\Gamma_{i}}u=u_{|\Gamma_{i}}$, one has $\operatorname{Tr}_{\Gamma_{1}}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma_{1})\right)$. When $\overline{\Gamma_{0}}\cap\overline{\Gamma_{1}}=\emptyset$ one also gets $\operatorname{Tr}_{\Gamma_{0}}\in{\mathcal{L}}\left(W^{m,p}(\Omega),W^{m-\frac{1}{p},p}(\Gamma_{0})\right)$.
Moreover, when $r\geq 2$ (i.e. $\Gamma$ is $C^{2}$), for all $p\in(1,\infty)$, $m\in\mathbb{N}$ such that $2\leq m\leq r$, and $u\in W^{m,p}(\Omega)$, denoting $\partial_{\nu}u=\sum_{i=1}^{N}\partial_{i}u_{|\Gamma}\,\nu^{i}$, where $\nu=(\nu^{1},\ldots,\nu^{N})$, and $\partial_{i}u_{|\Gamma_{i}}$ are taken in the trace sense sense, we get the normal derivative operator
@@
Clearly $\partial_{\nu}u$ (its restrictions to $\Gamma_{i}$ will be denoted by $\partial_{\nu}u_{|\Gamma_{i}}$) can be defined in such a trace sense only when $r\geq 2$ and $u\in W^{2,p}(\Omega)$. Hence, when $r=1$, we set the normal derivative in a distributional sense as follows.
@@
We recall, see [45, Lemma 1, p. 2147] (the quoted result trivially extends to $\Gamma$ of class $C^{1}$) that the space
@@
On the other hand, by the assumption $\mathcal{H}^{N-1}(\Gamma_{0})>0$, the norm $\|\cdot\|_{H^{1}}$ given by (9) is equivalent to ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{H^{1}}$, as it is well–known. We refer to [50, Lemma 1, p. 8] for an explicit proof. Hence in the paper, as stated in § 1, we shall endow the space $H^{1}$ with the norm $\|\cdot\|_{H^{1}}$ induced by the inner product $(\cdot,\cdot)_{H^{1}}$ defined in (9).
@@ 3. Analysis of problem (13)
This section is devoted to a preliminary analysis of problem (13). We start by making precise which types of solutions we are going to consider.
@@
    When assumption (R) holds, we say that $u\in H^{2}$ is a strong solution of (13) provided
@@
Essentially as in [47], it is useful to deal with weak solutions of (13) in a more abstract sense. By [47, Lemma 2.1], trivially extended to the complex case, the embedding $H^{1}\hookrightarrow H^{0}$ is dense. In the sequel we shal identify $L^{2}(\Omega)$ and $L^{2}(\Gamma_{1})$ with their duals $[L^{2}(\Omega)]^{\prime}$ and $[L^{2}(\Gamma_{1})]^{\prime}$, coherently with identifications usually made in the distribution sense. We shall also identify $L^{2}(\Omega)$ and $L^{2}(\Gamma_{1})$ with their isometric copies $L^{2}(\Omega)\times\{0\}$ and $\{0\}\times L^{2}(\Gamma_{1})$ contained in $H^{0}=L^{2}(\Omega)\times L^{2}(\Gamma_{1})$. Consequently, we shall identify $H^{0}$ with its dual $(H^{0})^{\prime}$, according to the identity
@@
We then introduce the chain of dense embeddings, or Gel’fand triple,
@@
(41) $\langle A_{1}u,v\rangle_{H^{1}}=(u,v)_{H^{1}}=\int_{\Omega}\nabla u\nabla v+\int_{\Gamma_{1}}(\nabla_{\Gamma}u,\nabla_{\Gamma}\overline{v})_{\Gamma},$
and its part $A:D(A)\subset H^{1}\subset H^{0}\to H^{0}$ given by
@@
were (39) was used. By a quick comparison between (37) and (41)–(42) one gets the following result.
@@
Combining it with (43) we get (37), that is $u$ is a weak solution of (13). By Lemma 8 thus $u\in D(A)$. Since all $u\in H^{2}$ are strong solutions of (13) for appropriate $(f,g)\in H^{0}$, we also get $H^{2}\subseteq D(A)$.
@@
Before stating our next result we remark that, since the operator $A$ is trivially closed, we can endow $D(A)$ with the graph norm $\|\cdot\|_{D(A)}=\|\cdot\|_{H^{1}}+\|A(\cdot)\|_{H^{0}}$, obtaining a Hilbert space. The following result shows that problem (13) is well–posed and $A$ is an isomorphism.
@@
    If $f\in L^{p}(\Omega)$ and $g\in L^{p}(\Gamma_{1})$ for $p\in[2,\infty)$, then $u\in W^{2,p}$. Moreover, there is a positive constant $c_{2}=c_{2}(\Omega,\Gamma_{1},p)$ such that
@@
    (46) $\|u\|_{H^{m}}\leq c_{3}\left(\|f\|_{H^{m-2}(\Omega)}+\|g\|_{H^{m-2}(\Gamma_{1})}\right)$
@@
We start by proving the first assertion in ii). We argue by induction on $m$. When $m=2$, by Lemma 9 we have $u\in H^{2}$. Now we suppose by the induction hypotesis that $m\geq 3$, the assertion holds for $m-1$, $f\in H^{m-2}(\Omega)$ and $g\in H^{m-2}(\Gamma_{1})$. We claim that $u\in H^{m}$. By the induction hypotesis we have $u\in H^{m-1}$. Consequently, by (33), $\partial_{\nu}u_{|\Gamma_{1}}\in H^{m-5/2}(\Gamma_{1})$. Since, by Lemma 9, we have $\partial_{\nu}u_{|\Gamma_{1}}-\Delta_{\Gamma_{1}}u=g$ on $\Gamma_{1}$, by Lemma 6 we get $u_{|\Gamma_{1}}\in H^{m-1/2}(\Gamma_{1})$. By the same Lemma we also have $-\Delta u=f$ in $\Omega$, so by elliptic higher regularity results (see [21, Chapter 2, Theorem 2.5.1.1, p. 128]) we get $u\in H^{m}(\Omega)$. A further application of (33) then yields $\partial_{\nu}u_{|\Gamma_{1}}\in H^{m-3/2}(\Gamma_{1})$. Since $\partial_{\nu}u_{|\Gamma_{1}}-\Delta_{\Gamma_{1}}u=g$ on $\Gamma_{1}$, again by Lemma 6 we get $u_{|\Gamma_{1}}\in H^{m}(\Gamma_{1})$, proving our claim.
@@
It is then clear that, going on in this way, the following alternative occurs: either after finitely many iterations we get $\partial_{\nu}u_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$, so the proof of our claim can be completed as when $N=2$, or we can go on indefinitely. In this case the recursive formula $p_{0}=2$, $p_{n+1}=\frac{(N-1)p_{n}}{N-p_{n}}$, defines a strictly increasing sequence $(p_{n})_{n}$ such that $1<p_{n}<\min\{p,N\}$ for all $n\in\mathbb{N}$. Let $l:=\lim_{n}p_{n}\in(1,\min\{p,N\}]$. Passing to the limit in the recursive formula we get that $l=N$ leads to $l=\infty$, a contradiction, so $l<N$. Passing to the limit then $l=\frac{(N-1)l}{N-l}$, that is $l=1$, a contradiction. Hence after finitely many iterations we get $\partial_{\nu}u_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ and conclude the proof of our claim. To prove (45) and conclude the proof one then applies the same argument already used to prove (46). ∎
@@ 4. Proofs
This section is devoted to prove our main results, stated in Section 1.
@@
We are then going to apply to $B$ the standard Spectral Decomposition Theorem for compact and self adjoint operators, see for example [6] or [14]. The spectrum $\sigma(B)$ of $B$ then consists of the union of $\{0\}$ and of its point spectrum $\sigma_{p}(B)\subset\mathbb{R}$, which is at most countable. Moreover, since $A$ is bijective, $0\not\in\sigma_{p}(B)$ and, by the Fredholm Alternative, $\text{ker}(B-\mu I)$ is finite–dimensional for all $\mu\in\sigma_{p}(B)$. Moreover $H^{0}$, being separable and infinite dimensional, admits a Hilbert basis of eigenvectors of $B$. Hence $\sigma_{p}(B)=\{\mu_{n},n\in\mathbb{N}\}$, where the sequence $(\mu_{n})_{n}$ is injective and $\mu_{n}\to 0$ as $n\to\infty$. Consequently $\Lambda=\{\lambda_{n},n\in\mathbb{N}\}$, where $\lambda_{n}=1/\mu_{n}$. Let us denote $H_{n}=\text{ker}(B-\mu_{n}I)$. Since, for each $u\in H_{n}$, we have $\lambda_{n}=\|u\|_{H^{1}}^{2}$, it follows that $\Lambda\subset(0,\infty)$. Consequently we can re–arrange the $\lambda_{n}$’s, according to their finite multiplicities, as in (11).
@@
At first we prove i), by combining Theorem 11–i) with a bootstrap argument. More in detail, by Lemma 9, $u_{n}\in H^{2}$ for each $n\in\mathbb{N}$, that is $u_{n}\in H^{2}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in H^{2}(\Gamma_{1})$. We now distinguish between the cases $N\leq 4$ and $N\geq 5$. In the first one, by applying Morrey’s Theorem when $N<4$ and the limiting case of Sobolev Embedding Theorem when $N=4$, we get $u_{n}\in L^{p}(\Omega)$ for all $p\in[2,\infty)$. By applying the same results on $\Gamma_{1}$ we get that ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ for all $p\in[2,\infty)$ provided $N\leq 5$, so a – fortiori when $N\leq 4$. By applying Theorem 11–i) we then get $u_{n}\in W^{2,p}$ for all $p\in[2,\infty)$. When $N\geq 5$ we further distinguish between the case $5\leq N\leq 8$ and $N\geq 9$. In the first one, by Sobolev Embedding Theorem, we have $u_{n}\in L^{q_{1}}(\Omega)$, where $\frac{1}{q_{1}}=\frac{1}{2}-\frac{2}{N}$, and as above one also gets that $u_{n}\in L^{q_{1}}(\Gamma_{1})$. Hence, by Theorem 11–i), one has $u_{n}\in W^{2,q_{1}}$. Since $\frac{1}{q_{1}}-\frac{2}{N}=\frac{1}{2}-\frac{4}{N}\leq 0$, by applying Morrey’s Theorem when $N<8$ and the limiting case of Sobolev Embedding Theorem when $N=8$, we get $u_{n}\in L^{p}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ for all $p\in[2,\infty)$, so $u_{n}\in W^{2,p}$ for all $p\in[2,\infty)$ as in the previous case. It is then clear how to handle the remaining case $N\geq 9$, proving after finitely many iterations that $u_{n}\in L^{p}(\Omega)$ and ${u_{n}}_{|\Gamma_{1}}\in L^{p}(\Gamma_{1})$ , so $u_{n}\in W^{2,p}$, for all $p\in[2,\infty)$. The conclusion $u_{n}\in C^{1}(\overline{\Omega})$ then follows from Morrey’s Theorem.
To prove ii) we remark that, since $u_{n}\in W^{2,p}$, for all $p\in[2,\infty)$, using (33) we have $\partial_{\nu}{u_{n}}_{|\Gamma_{1}}\in W^{1-\frac{1}{p},p}(\Gamma_{1})$ for all $p\in[2,\infty)$. We also remark that $W^{1-\frac{1}{p},p}(\Gamma_{1})$ coincides with the Besov space $B^{1-\frac{1}{p},p}_{p}(\Gamma_{1})$, as proved in [44, pp. 189–190]. By Morrey’s Theorem for Besov spaces, see [28, Chapter 17, Theorem 17.52, p. 562], taking $p>N$, we have $\partial_{\nu}{u_{n}}_{|\Gamma_{1}}\in C^{0,1-\frac{N}{p}}(\Gamma_{1})$. Moreover, since we also have ${u_{n}}_{|\Gamma_{1}}\in C^{1}(\Gamma_{1})$, by (12) and Lemma 9, $-\Delta_{\Gamma_{1}}u_{n}=f_{n}\in C^{0,1-\frac{N}{p}}(\Gamma_{1})$. Being $\Gamma_{1}$ compact, we can use the standard localization technique together with the classical Schauder Estimates, see [6, Theorem 9.33] or [19, Chapter 6], to recognize that in local coordinates $u_{n}$ is of class $C^{2,1-\frac{N}{p}}$, so $u_{n}\in C^{2}(\Gamma_{1})$, proving ii).
@@
Moreover, since $\Gamma$ is $C^{2}$, it is also $C^{1,1}$, so $\Omega$ satisfies the interior ball condition, that is to say there is an open ball $B\subseteq\Omega$ such that $x_{0}\in\partial B$. See [29, Corollary 2,p. 550] or [3, Theorem 1.0.9,p. 7]. Now, since $u\in C^{\infty}(B)\cap C^{1}(\overline{B})$, $-\Delta u=\lambda_{1}u\geq 0$ in $B$, and $u(x)>u(x_{0})=0$ for all $x\in B$, we can apply the classical Höpf Lemma, see [17, Chapter 6,p. 330], to conclude that $\partial_{\nu}u(x_{0})<0$, contradicting (60) and completing the proof. ∎